OpenStudy (anonymous):

Let f(x) = x2 – 2x + 2 and g(x) = x – 3. Find f(x) • g(x).

OpenStudy (anonymous):

I assume this means the composition, also f(g(x))? This means that everywhere you see an "x" in f(x), plug in g(x): f(g(x)) = (g(x))^2 - 2(g(x)) + 2 = (x - 3)^2 - 2(x - 3) + 2 = x^2 - 6x + 9 -2x + 6 + 2 = x^2 - 8x + 17 Unless you mean something else..

OpenStudy (anonymous):

yes i mean something else

OpenStudy (radar):

Is x^3-5x^2+8x-6 the desired results?

OpenStudy (anonymous):

yes thank you1!!!